OpenStudy (anonymous):

Choose the correct simplification of the expression 4/y^-3

OpenStudy (anonymous):

@rajathsbhat @AriPotta

OpenStudy (anonymous):

@uri

OpenStudy (raden):

use the property of exponent : 1/a^-m = a^m

OpenStudy (raden):

so, 4/y^-3 = 4 * 1/y^-3 = 4 * y^3

OpenStudy (anonymous):

\[4y ^{3}\]

OpenStudy (raden):

yep

OpenStudy (anonymous):

\[\left(\begin{matrix}2b \\ c\end{matrix}\right) ^{3}\]

OpenStudy (anonymous):

the answer would be 8b^3 / c^3 ?

OpenStudy (anonymous):

\[\frac{ 8b ^{3} }{ c ^{3} }\]

OpenStudy (raden):

if in bracket was a fraction, yes u are right :)

OpenStudy (anonymous):

ok :) im gonna put my other questions in other chats so u get more medals :)